OpenStudy (anonymous):

What is the third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3? f(x) = 2x3 + 12x2 – 22x – 12 f(x) = 2x3 – 12x2 + 22x – 12 f(x) = 4x3 – 24x2 + 22x – 24 f(x) = 4x3 + 24x2 – 44x – 24 What is the third–degree polynomial function such that f(0) = –12 and whose zeros are 1, 2, and 3? f(x) = 2x3 + 12x2 – 22x – 12 f(x) = 2x3 – 12x2 + 22x – 12 f(x) = 4x3 – 24x2 + 22x – 24 f(x) = 4x3 + 24x2 – 44x – 24 @Mathematics

OpenStudy (anonymous):

Some help?

OpenStudy (lalaly):

the factored form of a third degree polynomial looks like f(x) = k(x-a)(x-b)(x-c) f(0)=-12 that means when x =0 f(x)=-12 and a=1, b=2, c=3 substituting in equation 1 f(0)= k(0-1)(0-2)(0-3) f(0) = -6k therefore -12=-6k k=2 so the polynomial is f(x)=2(x-1)(x-2)(x-3)

OpenStudy (lalaly):

expand it lol

OpenStudy (lalaly):

The second choice is the right one btw:)

OpenStudy (anonymous):

haha k, how did you get -6?

OpenStudy (lalaly):

u know that f(0) = -12 right?

OpenStudy (anonymous):

yes

OpenStudy (lalaly):

and f(x)=k(x-a)(x-b)(x-c) so if u substitute x=0 here f(0)=k(0-a)(0-b)(0-c) f(0) = k(-a)(-b)(-c) a, b and c are the zeros of the polynomial so a=1 and b=2 and c=3 substituting these values in f(0) u get f(0) = k(-1)(-2)(-3) f(0)= k * -6 and we know that f(0)=-12 so -6k=-12

OpenStudy (anonymous):

Oh is this a theorem? or no?

OpenStudy (lalaly):

all third degree polynomials look like f(x)=k(x-a)(x-b)(x-c) where k is any constant and a,b,c are zeros of polynomial

OpenStudy (anonymous):

Oh ok, man I have to find that in my lessons, I dont even know if it mentioned it haha.

OpenStudy (lalaly):

hehe, it can be found using other methods, this is the easiest

OpenStudy (anonymous):

Oh would you mind telling me the other methods? sorry for all the questions haha I just want to have it down packed